OpenStudy (anonymous):

7y=y-42??? how do i solve the equation?

OpenStudy (anonymous):

6y = -42 y=-7

OpenStudy (anonymous):

7y=y-42 6y=-42 y=-42/6 y=-7

OpenStudy (anonymous):

how did yu get 6y?? @Silentempest